if 2 is subtracted from a number and this difference is tripled, the result is 6 more than the number. finf the nunber

Answer by ankor@dixie-net.com(22740)

If 2 is subtracted from a number and this difference is tripled,
the result is 6 more than the number. find the number
:
Let x = "a number"
:
3(x - 2) = x + 6
3x - 6 = x + 6
3x - x = 6 + 6
2x = 12
x = 6 is the number
: